# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

### v1

@@ Abstract
Keywords: transfer learning; high-dimensional quantile regression; distribution shift; transferable set; orthogonal debiasing
@@ 1 Introduction
Previous experiences can offer valuable insights for learning new tasks. Transfer learning, the technique of improving the performance of target learners on target domains by leveraging knowledge from different but related source domains (Zhuang et al.,, 2021), has shown significant success across various application scenarios. These applications range from machine learning problems such as natural language processing (Ruder et al.,, 2019) and computer vision (Shao et al.,, 2015), to science problems such as protein representation (Fenoy et al.,, 2022) and drug discovery (Turki et al.,, 2017), owing to the increasing richness of data sources. However, our understanding of transfer learning from a statistical perspective remains incomplete, with the fundamental challenge being the detection and utilization of valuable information within heterogeneous source studies.
In this paper, we focus on the transfer learning for quantile regression in the context of high-dimensional setting. Quantile regression (QR) models the impact of covariates on the conditional response distribution and has achieved great success in many applications. For QR in various areas, see Koenker, (2017) for a detailed review. One of QR’s most appealing features is its ability to describe the impact of covariates not only on the mean but also on the tails of the response distribution with robustness to heteroscedastic errors. Given potential non-negligible divergence between target and source distributions, QR can provide a more comprehensive assessment of what can be transferred from the source studies. Moreover, we focus on the regime of high-dimensional models, since the error bounds of high-dimensional estimators are heavily limited by the sample size. This motivates us to explore multiple external datasets to improve the performance of high-dimensional estimators on the target task.
To set up the framework, suppose we have a target dataset $((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}$ and the parameter of interest is the quantile of the response $y_{i}^{(0)}$ conditional on covariates $\boldsymbol{x}_{i}^{(0)}\in\mathbb{R}^{p}$ at a given quantile level $\tau\in(0,1)$, denoted by $\boldsymbol{\beta^{*}}$ (sometimes we use $\boldsymbol{w}^{(0)}$ for unity). Also, we have $K$ independent source datasets $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k=1}^{K}$, where the corresponding quantile coefficients are denoted by $\boldsymbol{w}^{(k)},k=1,\dots,K$. Denote the positive covariance matrix $\boldsymbol{\Sigma}^{(k)}:=\mathbb{E}[\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]$ for each study. In the high-dimensional setting, the number of covariates $p$ is very large, possibly larger than the overall sample size $N:=\sum_{k=0}^{K}n_{k}$. Assume a linear QR model, which is
@@
where each model residual $\epsilon_{i}^{(k)}$ satisfies $\mathbb{P}(\epsilon_{i}^{(k)}\leq 0|\boldsymbol{x}_{i}^{(k)})=\tau$ with conditional density function $f^{(k)}(\cdot|\boldsymbol{x}_{i}^{(k)})$. Assume the target parameter $\boldsymbol{\beta^{*}}$ is sparse, where the number of non-zero coefficients satisfies $\|\boldsymbol{\beta^{*}}\|_{0}=s\ll\min\{n_{0},p\}$. Also, we define the contrast vector $\boldsymbol{\delta}^{(k)}:=\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}$ (obviously $\boldsymbol{\delta}^{(0)}=0$). Our goal is to estimate the target parameter $\boldsymbol{\beta^{*}}$, with the help of abundant external data.
Three types of distribution shift. For the QR model (1) specified on all studies, there may exist three types of distribution shifts between the target and source domains:
  1. 1.
    parameter shift: $\boldsymbol{\beta^{*}}\neq\boldsymbol{w}^{(k)}$, i.e., the model coefficients differ.
  2. 2.
    residual shift: $P(\epsilon_{i}^{(0)}|\boldsymbol{x}_{i}^{(0)})\neq P(\epsilon_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})$, i.e., the conditional distributions of model residuals differ.
  3. 3.
    covariate shift: $P(\boldsymbol{x}_{i}^{(0)})\neq P(\boldsymbol{x}_{i}^{(k)})$, i.e., the marginal covariate distributions differ.
Both parameter shift and residual shift can be seen as the posterior/label shift, meaning that the conditional distribution of responses differs between $P(y_{i}^{(0)}|\boldsymbol{x}_{i}^{(0)})$ and $P(y_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})$. Many examples have revealed that transferring source studies under distribution shift may pose challenges to the final performance and lead to the negative transfer (Wang et al.,, 2019; Zhang et al.,, 2022). Hence, the core issue of this paper is:
@@
Methodologically, transfer learning falls within the field of multi-source data analysis, a prominent topic in contemporary statistics. Numerous approaches have been developed for multi-source settings, varying by task type, data category, and other factors. These include federated learning (McMahan et al.,, 2017), distributed learning (Jordan et al.,, 2019; Fan et al.,, 2023), fusion learning (Shen et al.,, 2020; Li and Luedtke,, 2023), multi-task learning (Duan and Wang,, 2023; Knight and Duan,, 2023), and transfer learning (Pan and Yang,, 2009; Zhuang et al.,, 2021), each with its unique characteristics and structures. For transfer learning, it aims to leverage information from source domains to improve the performance of the target task, with the core challenge being how to extract valuable source information while addressing data heterogeneity. Transfer learning has been studied across a variety of contexts, including linear regression (Li et al.,, 2022), generalized linear regression (Tian and Feng,, 2023), nonparametric classification (Cai and Wei,, 2021), nonparametric regression (Cai and Pu,, 2024), graphical models (Li et al.,, 2023), and causal inference (Wei et al.,, 2023).
The main focus of this paper lies on the transfer learning for high-dimensional linear models, where the goal is estimation and inference for the target parameter. For parameter estimation, a seminal and widely adopted approach is the two-step framework introduced in Li et al., (2022). In this method, the oracle transferable set $A_{h}:=\{\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h\}$ is defined, where source studies with a small contrast of parameters are considered informative. The two-step procedure first combines the data from $A_{h}$ to obtain a pooled estimate, then debiases this estimate using the target data. Tian and Feng, (2023) then extended this approach to generalized linear regression, providing a consistent source detection algorithm based on cross-validated loss comparison. Huang et al., (2023) and Zhang and Zhu, (2023) further developed direct extensions of the two-step framework and source detection procedure for high-dimensional quantile regression and smoothed quantile regression, respectively. Among the limited literature for statistical inference with knowledge transfer, Tian and Feng, (2023), Li et al., (2024), and Huang et al., (2023) applied the one-step correction technique (Zhang and Zhang,, 2014; van de Geer et al.,, 2014) to debias the previous transfer estimate on the target study.
As defined by the transferable set $A_{h}$, the two-step approach only addresses parameter shift, leaving it susceptible to covariate and residual shifts across multiple sources. However, studies in machine learning have shown that it is essential to account for both covariate and label shifts in heterogeneous data settings, such as image recognition Park et al., (2023) and disease diagnosis Chen et al., (2022). Additionally, due to these shifts, prior debiased transfer estimators achieved only $\sqrt{n_{0}}$-normality, providing no improvement over the original target baseline. To reduce the impact of covariate shift, Li et al., (2024) proposed jointly estimating the $\boldsymbol{w}^{(k)}$ of each study using $\ell_{1}$-minimization, Zhao et al., (2023) utilized a residual importance-weighted algorithm to make better use of source observations, and He et al., (2024) added a fused-regularizer to the pooling step of the two-step approach. While these works addressed covariate shift and introduced remedies, none of them account for residual shift. However, several studies suggest that addressing complex label shift issues requires a distributional approach, such as density estimation (Lipton et al.,, 2018; Garg et al.,, 2020; Lee et al.,, 2024), rather than simply assuming a linear relationship and comparing coefficients. In our context of quantile regression, the distributional information around each quantile differs across sources, making it essential to handle residual shift to accurately capture the information at each quantile of the target response. To demonstrate how severe residual shift can lead to negative transfer, we provide a motivating example through simulations of existing transfer methods.
The motivating example follows a similar setup to the simulation in Zhang and Zhu, (2023). We set $p=500,s=10,n_{0}=200,\boldsymbol{\beta^{*}}=(\mathbf{0.5}_{s},\mathbf{0}_{p-s})^{\top},\boldsymbol{x}_{i}^{(0)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$ with $\boldsymbol{\Sigma}_{\boldsymbol{x}}=(0.7^{|i-j|})_{1\leq i,j\leq p}$. We consider only one source study with sample size $n_{1}$ varying from 100 to 500, and we assume no parameter or covariate shift, i.e., $\boldsymbol{\delta}^{(1)}=0,\boldsymbol{x}_{i}^{(1)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$. The key change lies in the residual distributions, where $\epsilon_{i}^{(0)}$ is standard normal, but $\epsilon_{i}^{(1)}$ follows one of the four types of distributions: 1) standard normal: $\mathcal{N}(0,1)$; 2) Cauchy distribution: $\mathcal{C}(0,3)$; 3) mixed Gaussian: $z\mathcal{N}(-3,0.5)+(1-z)\mathcal{N}(3,0.5)$, $z\sim\mathcal{B}ernoulli(\tau)$; 4) noisy normal: $\mathcal{N}(0,5^{2})$. All residuals are generated with necessary translational adjustments for the $\tau$-th quantile at zero.
It is worth noting that in this setting, the source study is included in the oracle transferable set by the previous definition (Li et al.,, 2022; Tian and Feng,, 2023). We compare four methods for estimating the target parameter $\boldsymbol{\beta^{*}}$ at $0.2$-th quantile, which are: 1) L1-QR: $\ell_{1}$-penalized quantile regression (Belloni and Chernozhukov,, 2011) on the target data; 2) Pooling: $\ell_{1}$-penalized quantile regression on the entire dataset; 3) TransQR_2step: two-step transfer for quantile regression (Huang et al.,, 2023); 4) TransSQR_2step: two-step transfer for smoothed quantile regression (Zhang and Zhu,, 2023). We repeat these methods 50 times and report the average of $\ell_{2}$-estimation errors in Figure 1.1.
Figure 1.1: Average $\ell_{2}$-estimation errors of four methods under residual shift
@@
As shown in Figure 1.1, it is clear that, apart from the normal setting which demonstrates the anticipated transfer benefit, all other settings reveal negative transfer effects. Pooling, TransQR_2step, and TransSQR_2step perform significantly worse than the baseline L1-QR on the target data, strongly suggesting that assessing information transferability based solely on parameter similarity is insufficient in cases involving residual shift. This observation raises an interesting question: how can we effectively measure and control residual shift? Note that the adverse impact of residual shift fundamentally arises from information loss around the target quantile. So, what constitutes a meaningful representation of local distributional information? The answer is density. Motivated by this, we propose a novel transferable set:
@@
$\mathcal{C}_{h}=\{k:1\leq k\leq K,\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\mathbb{E}[f^{(0)}(0|\boldsymbol{x}_{i}^{(0)})]\leq h_{2}\mathbb{E}[f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})]\},$
@@
where $h_{1},h_{2}$ represent oracle yet unknown transferability levels. This set is designed to filter out source studies characterized by large parameter discrepancies and small densities at the $\tau$-th quantile, thus enabling the simultaneous management of parameter and residual shifts. We will further elaborate on the rationale behind the construction in Section 2. Using this transferable set, we develop a whole transfer learning framework to address parameter, covariate, and residual shifts concurrently within the high-dimensional quantile regression domain. Our contributions are:
  • We introduce the concept of residual shift in the multi-source setting and define a new transferable set to handle both parameter and residual shifts, offering deeper insights into managing distribution shift under knowledge transfer. For parameter estimation, we extend the constrained $\ell_{1}$-minimization algorithm (Li et al.,, 2024) to quantile regression, which remains robust in the presence of covariate shift. Furthermore, we develop a corresponding source detection procedure and integrate it into a unified one-shot estimation algorithm.
  • In theory, we derive non-asymptotic $\ell_{1}/\ell_{2}$-error bounds for our transfer estimators, broadening existing results to account for multiple types of distribution shifts, not just parameter shift. The resulting convergence rate surpasses that of single-task error bounds (Belloni and Chernozhukov,, 2011; Wang and He,, 2024), underscoring the advantage of our approach, especially when divergent sources are involved. The detection consistency of our source screening procedure is also established.
  • A debiased transfer approach based on Neyman’s orthogonality (Belloni et al.,, 2019) is proposed for statistical inference on quantile coefficients with knowledge transfer. Our method avoids the estimation of high-dimensional precision matrices and allows non-identical data distributions, which helps to alleviate the trouble by covariate and residual shift under knowledge transfer. To the best of our knowledge, our method is the first to debias on informative sources to enhance inference efficiency. By integrating valuable source studies, our approach achieves sharper asymptotic rate than previous results, leading to more efficient debiased estimates and shorter confidence intervals.
@@
The rest of the article is organized as follows. In Section 2, we propose our transfer learning approach under distribution shift. We will introduce our transferable set and transfer framework. Non-asymptotic error bounds are revealed subsequently. Section 3 describes the source detection procedure, accompanied by consistency guarantees. In Section 4, we present the orthogonal debiased approach for statistical inference with knowledge transfer, along with the corresponding asymptotic theory. Section 5 reports simulation results on parameter estimation and statistical inference, highlighting the advantages of our methods. Finally, in Section 6 we apply our methods to the Genotype-Tissue Expression (GTEx) dataset, predicting brain tissue gene expression levels by leveraging information from other tissues.
@@ 2 Transfer Learning under Distribution Shift
We begin this section with some necessary notations. For two sequences of positive numbers $\left\{a_{n}\right\}$ and $\left\{b_{n}\right\}$, we write $a_{n}\lesssim b_{n}$ if $a_{n}\leq cb_{n}$ for some universal constant $c\in(0,\infty)$, and $a_{n}\gtrsim b_{n}$ if $a_{n}\geq c^{\prime}b_{n}$ for some universal constant $c^{\prime}\in(0,\infty)$. We say $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. For any $q\in[0,\infty]$ and vector $\boldsymbol{x}=(x_{1},\cdots,x_{p})\in\mathbb{R}^{p}$, we write $\|\boldsymbol{x}\|_{q}$ for its $l_{q}$-norm. Let $a\wedge b$ and $a\vee b$ denote the minimum and maximum of $a$ and $b$, respectively. For a vector $\boldsymbol{v}\in\mathbb{R}^{p}$ and a subset $S\subseteq\{1,2,\dots,p\}$, we use $\boldsymbol{v}_{S}$ to denote the restriction of vector $\boldsymbol{v}$ to the index set $S$. The sub-Gaussian norm of a random variable $u\in\mathbb{R}$ is $\|u\|_{\psi_{2}}=\inf\{t>0:\mathbb{E}[\exp(u^{2}/t^{2})]\leq 2\}$ and the sub-Gaussian norm of a random vector $\boldsymbol{U}\in\mathbb{R}^{n}$ is $\|\boldsymbol{U}\|_{\psi_{2}}=\sup_{\|\boldsymbol{v}\|_{2}=1,\boldsymbol{v}\in\mathbb{R}^{n}}\|\langle\boldsymbol{U},\boldsymbol{v}\rangle\|_{\psi_{2}}$. The sub-exponential norm of a random variable $u\in\mathbb{R}$ is $\|u\|_{\psi_{1}}=\inf\{t>0:\mathbb{E}[|u|/t]\leq 2\}$ and the sub-exponential norm of a random vector $\boldsymbol{U}\in\mathbb{R}^{n}$ is $\|\boldsymbol{U}\|_{\psi_{1}}=\sup_{\|\boldsymbol{v}\|_{2}=1,\boldsymbol{v}\in\mathbb{R}^{n}}\|\langle\boldsymbol{U},\boldsymbol{v}\rangle\|_{\psi_{1}}$. The expectation symbol $\mathbb{E}[(\cdot)_{i}^{(k)}]:=\mathbb{E}_{(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)})}[(\cdot)_{i}^{(k)}]$ means obtaining expectation with the joint distribution $P(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)})$.

### v2

@@ Abstract
Keywords: Knowledge transfer; High-dimensional quantile regression; Distribution shift; Transferable set; Orthogonal debiasing.
@@ 1 Introduction
Previous experiences can offer valuable insights for learning new tasks. Transfer learning improves the performance of target learners by leveraging knowledge from different but related sources [zhuang2020comprehensive], has achieved success across various areas, including natural language processing [ruder2019transfer], protein representation [fenoy2022transfer], and drug discovery [turki2017transfer]. Despite its broad applications, statistical understanding of transfer learning remains incomplete, with the fundamental challenge being the detection and integration of valuable information from heterogeneous sources.
In this paper, we focus on the transfer learning for quantile regression (QR) [koenker2017quantile] in the context of a high-dimensional setting. One of QR’s most appealing features is its ability to describe the impact of covariates not only on the mean but also on the tails of the response distribution with robustness to heteroscedastic errors. Given potential non-negligible divergence between target and source distributions, QR can provide a more comprehensive assessment of what can be transferred from the source studies.
To set up the framework, suppose we have a target dataset $((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}$ and $K$ independent source datasets $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k=1}^{K}$. Consider the linear QR models
@@
where $\boldsymbol{w}^{(k)}\in\mathbb{R}^{p}$ is the quantile coefficient of the response $y_{i}^{(k)}$ conditional on covariates $\boldsymbol{x}_{i}^{(k)}\in\mathbb{R}^{p}$ at a given quantile level $\tau\in(0,1)$, and $\epsilon_{i}^{(k)}$ is the error term satisfying $\mathbb{P}(\epsilon_{i}^{(k)}\leq 0|\boldsymbol{x}_{i}^{(k)})=\tau$ with conditional density function $f^{(k)}(\cdot|\boldsymbol{x}_{i}^{(k)})$. The parameter of interest is $\boldsymbol{\beta^{*}}:=\boldsymbol{w}^{(0)}$ on the target data, which is assumed to be sparse, i.e., $\|\boldsymbol{\beta^{*}}\|_{0}=s\ll\min\{n_{0},p\}$. In the high-dimensional setting, the number of covariates $p$ is very large, possibly larger than the overall sample size $N:=\sum_{k=0}^{K}n_{k}$. Our goal is to estimate the target parameter $\boldsymbol{\beta^{*}}$, with the help of abundant external data.
Three types of distribution shift. For the QR model (1) specified in all studies, there may exist three types of distribution shift between the target and source domains, which are: 1) parameter shift: $\boldsymbol{\beta^{*}}\neq\boldsymbol{w}^{(k)}$, i.e., the model coefficients differ; 2) residual shift: $P(\epsilon_{i}^{(0)}|\boldsymbol{x}_{i}^{(0)})\neq P(\epsilon_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})$, i.e., the conditional distributions of model residuals differ; 3) covariate shift: $P(\boldsymbol{x}_{i}^{(0)})\neq P(\boldsymbol{x}_{i}^{(k)})$, i.e., the marginal covariate distributions differ. Both parameter and residual shift can be seen as label shift, meaning that the conditional label distributions $P(y_{i}^{(0)}|\boldsymbol{x}_{i}^{(0)})$ and $P(y_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})$ differ. Many examples have revealed that blindly transferring sources under distribution shift may pose challenges to the final performance and lead to the negative transfer [wang2019characterizing, zhang2022survey]. Hence, the core issue of this paper is:
@@
Closely related to transfer learning, numerous approaches have been developed for multi-source data analysis, including federated learning [mcmahan2017communication], distributed learning [jordan2018communication, fan2023communication], fusion learning [shen2020fusion, li2023efficient], and multi-task learning [duan2023adaptive, knight2024multi]. However, distinguished from multi-source settings where each dataset is equally important, transfer learning aims to leverage information from source domains to improve the performance of the target task. Statistical investigations into transfer learning have spanned diverse contexts, including linear regression [li2022transfer, tian2022transfer], nonparametric classification [cai2021transfer], nonparametric regression [cai2024transfer], graphical models [li2023transfer], and causal inference [wei2023transfer].
To address parameter shift, [li2022transfer] introduced the contrast vector $\boldsymbol{\delta}^{(k)}:=\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}$ and defined the transferable set $A_{h}:=\{\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h\}$, where source studies with a small contrast of parameters are considered informative. They proposed a two-step transfer procedure to first obtain an integrated estimate based on informative sources, then debias this estimate on the target data. The two-step framework is further extended to generalized linear regression [tian2022transfer] and quantile regression [zhang2022transfer, huang2023estimation]. In addition, several studies have investigated covariate shift and proposed remedies in high-dimensional linear regression [li2023estimation, zhao2023residual, he2024transfusion].
To the best of our knowledge, none have explicitly accounted for residual shift. However, many applications highlight the necessity of addressing residual shift in heterogeneous settings, such as in image recognition [park2023label] and disease diagnosis [chen2022estimating]. Here a motivating example is provided to illustrate the potential negative transfer caused by residual shift. We consider a similar setting to the simulation in [zhang2022transfer] with no parameter or covariate shift. Specifically, we assume only one source study with different residual distributions. Figure 1 compares the average $\ell_{2}$-errors of four methods for estimating the target parameter at $0.2$-th quantile. More details can be found in the supplementary materials.
Figure 1: Average $\ell_{2}$-errors at $0.2$-th quantile, where source residual distributions are: 1) Normal: $\mathcal{N}(0,1)$; 2) Cauchy: $\mathcal{C}(0,5)$; 3) Mixed: mixed $\mathcal{N}(-3,0.5)$ and $\mathcal{N}(3,0.5)$; 4) Noisy: $\mathcal{N}(0,5^{2})$.
@@
As shown in Figure 1, apart from the normal setting that demonstrates the anticipated effect, all other settings reveal significant negative transfer, strongly suggesting that assessing transferability based solely on parameter similarity is insufficient in cases involving residual shift. Notably, residual shift fundamentally reflects the information quality of a source study around the objective quantile, where this difference can be characterized by the local density and sample size. Accordingly, we introduce a novel transferable set that simultaneously accounts for parameter and residual shift,
@@
$\mathcal{C}_{h}=\left\{k:1\leq k\leq K,\quad\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\quad\frac{n_{0}\mathbb{E}[f^{(0)}(0|\boldsymbol{x}_{i}^{(0)})]}{n_{k}\mathbb{E}[f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})]}\leq h_{2}\right\},$
@@
where $h_{1},h_{2}$ represent oracle yet unknown transferability levels. This set is designed to filter out source studies with large parameter discrepancies or limited information at the $\tau$-th quantile. We further elaborate on the rationale behind this set in Section 2. With this transferable set, we develop a comprehensive transfer learning framework to address parameter/covariate/residual shift concurrently for high-dimensional quantile regression.
Our contributions are threefold. First, we introduce a new transferable set to handle both parameter and residual shift, offering deeper insights into managing distribution shift under knowledge transfer. For parameter estimation, we extend the constrained $\ell_{1}$-minimization [li2023estimation] to quantile regression, ensuring robustness in the presence of covariate shift. We further develop a corresponding source detection procedure and integrate it into a unified transfer estimation algorithm. Second, in theory, we establish non-asymptotic upper error bounds for our transfer estimators, broadening existing results to account for residual shift in addition to parameter shift. We also derive the minimax lower error bound to demonstrate the rationale of our approach, and provide the detection consistency of the source screening procedure. Third, a debiased transfer approach based on Neyman orthogonality [belloni2019valid] is proposed for statistical inference on quantile coefficients, which to our knowledge, is the first to debias on informative sources to enhance inference efficiency under distribution shift. By integrating valuable source studies, our approach achieves sharper asymptotic rates than existing results with $\sqrt{n_{0}}$-normality [tian2022transfer, huang2023estimation], leading to more efficient debiased estimates and shorter confidence intervals.
@@
The rest of the article is organized as follows. In Section 2, we introduce our transfer framework under distribution shift. Section 3 describes the source detection procedure. In Section 4, we present the orthogonal debiased approach for statistical inference with knowledge transfer. Simulation studies and an application to Genotype-Tissue Expression (GTEx) dataset are provided in Section 5 and 6, respectively.
@@ 2 Transfer Learning under Distribution Shift
We begin this section with some necessary notations. For two sequences of positive numbers $\left\{a_{n}\right\}$ and $\left\{b_{n}\right\}$, we write $a_{n}\lesssim b_{n}$ if $a_{n}\leq cb_{n}$ for some universal constant $c\in(0,\infty)$, and $a_{n}\gtrsim b_{n}$ if $a_{n}\geq c^{\prime}b_{n}$ for some universal constant $c^{\prime}\in(0,\infty)$. We say $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. For any $q\in[0,\infty]$ and vector $\boldsymbol{x}=(x_{1},\cdots,x_{p})\in\mathbb{R}^{p}$, we write $\|\boldsymbol{x}\|_{q}$ for its $l_{q}$-norm. Let $a\wedge b$ and $a\vee b$ denote the minimum and maximum of $a$ and $b$, respectively. For an index subset $S\subseteq\{1,2,\dots,p\}$, let $|S|$ denote its cardinality, and let $S^{c}$ denote its complement, i.e., $S^{c}=\{1,2,\dots,p\}\setminus S$. For a vector $\boldsymbol{v}\in\mathbb{R}^{p}$, we use $\boldsymbol{v}_{S}$ to denote the restriction of vector $\boldsymbol{v}$ to the index set $S$. Denote $\boldsymbol{e}_{j}=(\boldsymbol{0}_{j-1}^{\top},1,\boldsymbol{0}_{p-j}^{\top})^{\top}$ as the unit vector. The sub-Gaussian norm of a random variable $u\in\mathbb{R}$ is $\|u\|_{\psi_{2}}=\inf\{t>0:\mathbb{E}[\exp(u^{2}/t^{2})]\leq 2\}$ and the sub-Gaussian norm of a random vector $\boldsymbol{U}\in\mathbb{R}^{n}$ is $\|\boldsymbol{U}\|_{\psi_{2}}=\sup_{\|\boldsymbol{v}\|_{2}=1,\boldsymbol{v}\in\mathbb{R}^{n}}\|\langle\boldsymbol{U},\boldsymbol{v}\rangle\|_{\psi_{2}}$. The expectation symbol $\mathbb{E}[(\cdot)_{i}^{(k)}]:=\mathbb{E}_{(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)})}[(\cdot)_{i}^{(k)}]$ means obtaining expectation with the joint distribution $P(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)})$.